what is the center and length of a circle with the equation 81x^2 + 81y^2 = 1?
Answer by lwsshak3(11628)

what is the center and length of a circle with the equation 81x^2 + 81y^2 = 1?
**
Standard form of an equation for a circle: (x-h)^2+(y-k)^2=r^2, (h,k) being the (x,y) coordinates of the center, r=radius.
..
81x^2 + 81y^2 = 1
Divide by 81
x^2+y^2=1/81
center: (0,0)
radius: √(1/81)=1/9 (I assumed this what you meant by length of a circle)